# Causal Structure Learning in Hawkes Processes with Complex Latent Confounder Networks

### v1

@@ Appendix E Accounting for Self-Looped Observed Subprocesses under Latent Confounder Influence
Hawkes Process Data: We generate event sequences using the tick library [2], an efficient framework for simulating multivariate Hawkes processes. The excitation function is set as exponential kernel $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$, where $\beta$ is fixed at 1. $\alpha_{ij}$ is sampled uniformly from $[0.8,0.99]$ except for Case 1. Because of the cycles between $N_{2}$ and $N_{3}$ of Case 1, large $\alpha_{ij}$ may lead to nonstationarity. Thus, we sample $\alpha_{ij}$ uniformly from $[0.40,0.80]$ specifically for Case 1. To ensure stationarity and avoid explosive behavior, we verify the spectral radius of the integrated excitation matrix after generating $\alpha_{ij}$. To discretize the sequences for our method, we select the time bins of length 0.1 and consider 600 effective lag time bins as discretized lag variables for the calculation sub-covariance matrix. The sample size corresponds to the number of discrete data points.
Discrete-Time Series Data: To assess our method under ideal discrete-time conditions (i.e., exactly satisfying Theorem˜3.1), we generate data directly from Eq. (2). The excitation function is set as exponential kernel $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$. The coefficients $\alpha_{ij}$ and decay parameter $\beta$ are set as above. Similar to the Hawkes data, we verify the spectral radius to ensure stationarity. The noise terms are drawn from independent Gaussian distributions. We set the number of effective lag variables to 200. The sample size corresponds to the number of discrete data points.
Preprocessing and Rank Deficiency Testing: For each trial, we standardize the discretized data to ensure fair comparison. To test for rank deficiency, we use canonical correlation analysis (CCA) [52], following the procedure in [22].
Data Usage for Baselines: For Hawkes process-based methods (SHP [39], THP [7], and NPHC [1]), we use the raw Hawkes process data produced by the tick library. For rank-based methods designed for i.i.d. data with linear relations (Hier. Rank [22] and RLCD [15]), we use the discretized Hawkes process data.
We run all the experiments on a personal PC (CPU).
@@
Causal relationships among both latent and observed subprocesses are represented by an adjacency matrix, where each entry is either 1 or 0, indicating the presence or absence of a directed edge, respectively. Specifically, $\text{Adj}{\mathcal{G}}(i,j)=1$ denotes a directed edge from the $j$-th subprocess to the $i$-th subprocess, while $\text{Adj}{\mathcal{G}}(i,j)=0$ indicates no such edge.
@@
In practice, the indices of latent subprocesses in the estimated (summary) graph may not correspond to those in the ground truth. To address this, following Huang et al. [22], we permute the latent subprocess indices in the estimated graph and select the permutation that minimizes the difference from the true graph. When the number of estimated latent subprocesses is smaller than the true number, we add isolated latent nodes to balance the comparison. Conversely, if the estimate exceeds the true number, we select the subset that best matches the true latent subprocesses.
Additionally, since our inferred summary graph simplifies the underlying causal structure, by omitting intermediate latent subprocesses and redundant edges as formalized in our theorems and Definition˜3.4, we adjust the ground-truth adjacency matrix to this idealized representation before comparison. This ensures a fair evaluation of causal discovery.
For baselines designed for i.i.d. data with linear relations (i.e., Hier. Rank [22] and RLCD [15]), their output graphs capture relationships among discretized variables, rather than subprocesses. To enable fair comparison, we regard an edge $N_{1}\rightarrow N_{2}$ as correctly identified if more than half of the considered variables associated with $N_{1}$ have inferred edges to those of $N_{2}$.
Q.3 Additional Experimental Results
@@
We evaluate the sensitivity of our method to the choice of the discretization interval $\Delta$ with decay parameter $\beta=1$ in the exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$. As shown in Table˜1, when $\Delta$ is set to 0.01 or 0.05, our method achieves consistently high F1-scores across all cases, confirming that the discretized representation sufficiently preserves the temporal dynamics of the underlying Hawkes process. Even at $\Delta=0.1$, the performance remains stable. However, when $\Delta$ increases to 0.3, we observe a sharp drop in performance, highlighting that overly coarse discretization leads to significant loss of temporal resolution, impairing the estimation of causal structures. The result shows the need to choose a small bin width $\Delta$ relative to the typical support of the excitation function [29, 32].
Robustness to Violations of Rank Faithfulness
To test robustness under violations of rank faithfulness, we randomly select two edges in each synthetic graph and assign them identical coefficients $\alpha_{ij}$ for the exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$ in every run. This manipulation introduces potential linear dependencies in the cross-covariance matrix, which could challenge rank-based methods. As presented in Table˜2, despite the induced degeneracy, our method maintains strong performance, especially as the sample size increases. These results suggest that our approach is robust to moderate violations of rank faithfulness in practical scenarios.
@@
We further evaluate our method on a larger causal graph with 14 subprocesses, as shown in Fig. 9. Table˜3 reports the F1-scores averaged over ten runs. Despite the increased complexity, our method successfully recovers the underlying causal structure with high accuracy.
@@
We evaluate our method on a real-world cellular network dataset [39], which includes expert-validated ground-truth causal relationships. The dataset comprises 18 distinct alarm types and approximately 35,000 recorded alarm events collected over eight months from an operational telecommunication network. For our evaluation, we focus on a subgraph involving five alarm types (Alarms 1–4 and 8), where Alarm 8 is manually excluded and treated as a latent subprocess. Both Alarm 2 and Alarm 4 are observed effects of this latent subprocess, providing an opportunity to assess our method’s ability to infer latent confounders. The ground-truth causal graph is shown in Figure˜10.
@@
Compared to our inferred causal graph, the ground truth contains an additional edge from Alarm 2 to Alarm 4. However, as noted in Definition˜3.4, causal edges between observed effects of a latent confounder are permissible in our framework.
@@
During inference, using Propositions˜3.3 and 3.7, we correctly identify Alarms 1 and 2 as the parent causes of Alarm 3, and Alarms 2 and 4 as the parent causes of Alarm 1. However, the parent cause sets of Alarms 2 and 4 cannot be fully explained by the observed subprocesses alone. This necessitates the existence of a latent confounder influencing both, leading to the successful identification of Alarm 8 as a latent subprocess.
@@
Figure 10: Ground-truth causal subgraph from the metropolitan cellular network dataset. Alarm 8 is treated as a latent subprocess.
Dataset Description
@@
The dataset records 34,838 alarm events from a metropolitan cellular network [39], covering 18 alarm types and 55 devices. Each record contains:
  • Alarm ID: The alarm type triggered (18 types).
  • Device ID: The device that generated the alarm (55 unique devices).
  • Start Timestamp: The time when the alarm was triggered.
  • End Timestamp: The time when the alarm was resolved.

### v2

@@ Appendix E Accounting for Self-Looped Observed Subprocesses under Latent Confounder Influence
Hawkes Process Data: We generate event sequences using the tick library [Bacry et al., 2017], an efficient framework for simulating multivariate Hawkes processes. The excitation function is set as exponential kernel $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$, where $\beta$ is fixed at 1. $\alpha_{ij}$ is sampled uniformly from $[0.8,0.99]$ except for Case 1. Because of the cycles between $N_{2}$ and $N_{3}$ of Case 1, large $\alpha_{ij}$ may lead to nonstationarity. Thus, we sample $\alpha_{ij}$ uniformly from $[0.40,0.80]$ specifically for Case 1. To ensure stationarity and avoid explosive behavior, we verify the spectral radius of the integrated excitation matrix after generating $\alpha_{ij}$. To discretize the sequences for our method, we select the time bins of length 0.1 and consider 600 effective lag time bins as discretized lagged variables for the calculation sub-covariance matrix. The sample size corresponds to the number of discrete data points.
Discrete-Time Series Data: To assess our method under ideal discrete-time conditions (i.e., exactly satisfying Theorem˜4.1), we generate data directly from Eq. (2). The excitation function is set as exponential kernel $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$. The coefficients $\alpha_{ij}$ and decay parameter $\beta$ are set as above. Similar to the Hawkes data, we verify the spectral radius to ensure stationarity. The noise terms are drawn from independent Gaussian distributions. We set the number of effective lagged variables to 200. The sample size corresponds to the number of discrete data points.
Preprocessing and Rank Deficiency Testing: For each trial, we standardize the discretized data to ensure fair comparison. To test for rank deficiency, we use canonical correlation analysis (CCA) [w Anderson, 1974], following the procedure in [Huang et al., 2022]. We use the grid search to find the best rank test threshold. We also conduct a empirical sensitivity analysis for test threshold. The result is in Table˜2. A threshold of 0.10 provides a good balance across multiple scenarios.
Data Usage for Baselines: For Hawkes process-based methods (SHP [Qiao et al., 2023], THP [Cai et al., 2022], and NPHC [Achab et al., 2018]), we use the raw Hawkes process data produced by the tick library. For rank-based methods designed for i.i.d. data with linear relations (Hier. Rank [Huang et al., 2022] and RLCD [Dong et al., 2023]) and time series baseline (LPCMCI [Gerhardus and Runge, 2020], we use the discretized Hawkes process data.
We run all the experiments on a personal CPU machine.
@@
Causal relationships among both latent and observed subprocesses are represented by an adjacency matrix, where each entry is either 1 or 0, indicating the presence or absence of a directed edge, respectively. Specifically, $\text{Adj}_{\mathcal{G}}(i,j)=1$ denotes a directed edge from the $j$-th subprocess to the $i$-th subprocess, while $\text{Adj}_{\mathcal{G}}(i,j)=0$ indicates no such edge.
@@
In practice, the indices of latent subprocesses in the estimated (summary) graph may not correspond to those in the ground truth. To address this, following Huang et al. Huang et al. [2022], we permute the latent subprocess indices in the estimated graph and select the permutation that minimizes the difference from the true graph. When the number of estimated latent subprocesses is smaller than the true number, we add isolated latent nodes to balance the comparison. Conversely, if the estimate exceeds the true number, we select the subset that best matches the true latent subprocesses.
Additionally, since our inferred summary graph simplifies the underlying causal structure, by omitting intermediate latent subprocesses and redundant edges as formalized in our theorems and Definition˜4.4, we adjust the ground-truth adjacency matrix to this idealized representation before comparison. This ensures a fair evaluation of causal discovery.
For baselines designed for i.i.d. data with linear relations (i.e., Hier. Rank [Huang et al., 2022] and RLCD [Dong et al., 2023]), their output graphs capture relationships among discretized variables, rather than subprocesses. To enable fair comparison, we regard an edge $N_{1}\rightarrow N_{2}$ as correctly identified if more than half of the considered variables associated with $N_{1}$ have inferred edges to those of $N_{2}$.
Q.3 Additional Synthetic Experiment Results
@@
We evaluate the sensitivity of our method to the choice of the discretization interval $\Delta$ with decay parameter $\beta=1$ in the exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$. As shown in Table˜1, when $\Delta$ is set to 0.01 or 0.05, our method achieves consistently high F1-scores across all cases, confirming that the discretized representation sufficiently preserves the temporal dynamics of the underlying Hawkes process. Even at $\Delta=0.1$, the performance remains stable. However, when $\Delta$ increases to 0.3, we observe a sharp drop in performance, highlighting that overly coarse discretization leads to significant loss of temporal resolution, impairing the estimation of causal structures. The result shows the need to choose a small bin width $\Delta$ relative to the typical support of the excitation function [Kirchner, 2016a].
@@
Sensitivity to Rank-Test Threshold
We evaluate the sensitivity of our method to the threshold $\tau$ used in the rank test (i.e., the cutoff deciding rank deficiency). We vary $\tau\in\{0.01,0.05,0.10,0.20\}$ and assess three representative cases. Each experiment uses 30k Hawkes samples generated by the tick library under an exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$ with $\beta=1$ and time interval $\Delta=0.1$; results are averaged over ten runs. As shown in Table˜2, in the fully observed setting (Case 1) precision remains $1.00$ while recall decreases as $\tau$ increases, whereas in latent settings (Cases 2–3) a moderately larger threshold improves precision because of the attenuation of causal influences through the latent subprocesses. Overall, a threshold of 0.10 provides a good balance across different scenarios.
Table 2: Sensitivity to the rank-test threshold $\tau$. Each entry is averaged over ten runs on 30k samples generated with an exponential kernel ($\beta=1$). Case 1–3 correspond to Figs. 1(b), 2(a), and 3(a), respectively. Overall, a threshold of 0.10 provides a good balance across different scenarios.
Precision Recall $F_{1}$
Threshold $\tau$ Case 1 Case 2 Case 3 Case 1 Case 2 Case 3 Case 1 Case 2 Case 3
0.01 1.00 0.42 0.57 0.80 0.53 0.50 0.88 0.47 0.53
0.05 1.00 0.62 0.62 0.64 0.73 0.54 0.77 0.67 0.57
0.10 1.00 0.66 0.72 0.60 0.75 0.65 0.74 0.71 0.68
0.20 1.00 0.76 0.68 0.47 0.85 0.63 0.62 0.80 0.65
Robustness to Violations of Rank Faithfulness
To test robustness under violations of rank faithfulness, we randomly select two edges in each synthetic graph and assign them identical coefficients $\alpha_{ij}$ for the exponential excitation function $\phi_{ij}(s)=\alpha_{ij}e^{-\beta s}$ in every run, with $\beta=1$ and time interval $\Delta=0.1$. This manipulation introduces potential linear dependencies in the cross-covariance matrix, which could challenge rank-based methods. As presented in Table˜3, despite the induced degeneracy, our method maintains strong performance, especially as the sample size increases. These results suggest that our approach is robust to moderate violations of rank faithfulness in practical scenarios.
@@
We further evaluate our method on a larger causal graph with 14 subprocesses, as shown in Fig. 9. Table˜4 reports the F1-scores averaged over ten runs. Despite the increased complexity, our method successfully recovers the underlying causal structure with high accuracy.
@@
Scalability and Runtime Profiling
We profile runtime on three representative synthetic graphs (Cases 1–3) and two real-world settings. All runs were executed on an AMD EPYC 9454 CPU. The first real-world setting follows our main paper: a five-alarm subgraph (Alarm_ids=0-3 with one latent Alarm_id=7) from device_id = 8. The second merges all devices into a single multivariate event sequence with all 18 alarms to gauge scaling with graph size. We observe that Case 1 is fastest as no latent confounders are present and Phase I suffices. Case 2 introduces latent confounders, requiring both phases in the first iteration and increasing runtime. Case 3 is slowest among synthetic cases because the latent confounder is itself caused by an observed subprocess, triggering an additional iteration to identify its observed parent. For real data, merging all devices markedly increases runtime as the sequence spans all 18 alarms and may deviate from a homogeneous Hawkes mechanism. A phase-wise complexity breakdown is provided in Appendix˜P, which offers further insight into the scalability of the algorithm.
Table 5: Runtime across synthetic and real-world settings.
Graph Type Runtime (s)
Case 1 227.80
Case 2 1036.01
Case 3 2603.95
Real Dataset (Alarm_ids=0--3, device_id = 8) 1364.71
Real Dataset (all devices merged; 18 alarms) 20914.29
@@
We evaluate our method on a real-world cellular network dataset [Qiao et al., 2023], which includes expert-validated ground-truth causal relationships. The dataset comprises 18 distinct alarm types and $\sim$35,000 recorded alarm events collected over eight months from an operational telecommunication network. This benchmark has been widely used in prior work (e.g., the PCIC 2021 causal discovery track and [Qiao et al., 2023]), where performance for many methods is available and top F1-scores are reported up to $\approx 0.6$.
@@
For our evaluation, we focus on a subgraph involving five alarm types (Alarm_ids=0-3 and 7) from device_id = 8, where Alarm_id=7 is manually excluded and treated as a latent subprocess. Both Alarm_id=1 and Alarm_id=3 are observed effects of this latent subprocess, providing an opportunity to assess our method’s ability to infer latent confounders. The ground-truth causal subgraph is shown in Figure˜10. Compared with our inferred causal graph, the ground truth contains an additional edge from Alarm_id=1 to Alarm_id=3. However, as noted in Definition˜4.4, causal edges between observed effects of a latent confounder are permissible in our framework.
@@
During inference, using Propositions˜4.3 and 4.7, we correctly identify Alarm_ids=0,1,3 as the parent causes of Alarm_id=2, and Alarm_ids=1,3 as the parent causes of Alarm_id=0. The parent-cause sets of Alarm_id=1 and Alarm_id=3 cannot be fully explained by the observed subprocesses alone. This necessitates the existence of a latent confounder influencing both, leading to the successful identification of Alarm_id=7 as a latent subprocess.
@@
Figure 10: Ground-truth causal subgraph from the metropolitan cellular network dataset. Alarm_id=7 is treated as a latent subprocess.
Baselines and Schemes.
We also quantitatively compare our method with representative Hawkes-based approaches (SHP [Qiao et al., 2023], THP [Cai et al., 2022], NPHC [Achab et al., 2018]), two rank-based latent-variable methods for i.i.d. data (Hier. Rank [Huang et al., 2022], RLCD [Dong et al., 2023]), and a time-series method designed for exogenous latents (LPCMCI [Gerhardus and Runge, 2020]). For fairness, all baselines are evaluated on the same sub-dataset (Alarm_ids=0--3 and 7 from device_id=8) as our method, with performance averaged over ten runs using parameter grid search.
Results on the Sub-dataset.
Our method achieves the best F1-score when the data conforms to a single multivariate Hawkes process (per-device setting). Table 6 reports the average F1-scores.
Table 6: F1-scores on the cellular network sub-dataset (Alarm_ids=0-3 and 7, device_id = 8) where Alarm_id=7 is manually excluded and treated as a latent subprocess; averages over 10 runs.
Algorithm F1-score
SHP 0.49
THP 0.48
NPHC 0.42
Hier. Rank 0.00
RLCD 0.39
LPCMCI 0.43
Ours 0.76
Merged-devices analysis.
For completeness, we also merge events from all 55 devices into a single multivariate sequence with all 18 alarm types and analyze it with our method. This setting violates the assumption that samples share the same generative mechanism (devices can be heterogeneous), and it yields a much lower F1-score (0.17). This illustrates why per-device analysis is more compatible with our assumptions, whereas merged-device data can confound structure learning.
Dataset description.
@@
The dataset records 34,838 alarm events from a metropolitan cellular network [Qiao et al., 2023], covering 18 alarm types and 55 devices. Each record contains:
  • Alarm ID: one of 18 alarm types,
  • Device ID: one of 55 devices,
  • Start Timestamp: time when the alarm was triggered,
  • End Timestamp: time when the alarm was resolved.